# Growth and characterization of the La3Ni2O7-δ thin films: dominant contribution of the orbital at ambient pressure

### v1

@@ Abstract
By using the pulsed-laser-ablation technique, we have successfully grown the La3Ni2O7-δ thin films with -axis orientation perpendicular to the film surface. X-ray diffraction shows that the (00l) peaks can be well indexed to the La3Ni2O7-δ phase. Resistive measurements show that the samples can be tuned from weak insulating to metallic behavior through adjusting the growth conditions. Surprisingly, all curves of in the temperature region of 2300 K do not show the anomalies corresponding to either the spin density wave or the charge density wave orders as seen in bulk samples. Hall effect measurements show a linear field dependence with the dominant hole charge carriers, but the Hall coefficient exhibits strong temperature dependence. The magnetoresistance above about 50 K is positive but very weak, indicating the absence of multiband effect. However, a negative magnetoresistance is observed at low temperatures, which shows the delocalization effect. Detailed analysis on the magnetoresistance suggests that the delocalization effect at low temperatures is due to the Kondo-like effect, rather than the Anderson weak localization. Our transport results suggest that, the electronic conduction is fulfilled by the orbital with holes as the dominant charge carriers, while the interaction through Hund’s coupling with the localized orbital plays an important role in the charge dynamics.
@@ II EXPERIMENTAL DETAILS
The La3Ni2O7-δ thin films were grown on single-crystal (00)-oriented LaAlO3 substrates by the technique of pulsed-laser deposition, with substrate temperature at and laser frequency of 2 Hz for all cases. Prior to the deposition, a polycrystalline La3Ni2O7-δ target was prepared by the sol-gel method [38]. During the growth, the laser spot area of 4.68 was fixed, and two laser fluence of 1.52 (labelled as SL) and 2.07 (labelled as SH) were applied. And the oxygen partial pressure during the growth was 2035 Pa. After growth, the samples were cooled down at a cooling rate of to under oxygen partial pressure =58 Pa and post-annealed in-situ at for about 148 h, followed by the same cooling procedure to room temperature. Then they were taken out for the subsequent measurements. XRD measurements were performed by using a Bruker D8 Advanced diffractometer with Cu radiation. The morphology characterization obtained by a scanning electron microscope (SEM) (Fig. 2(a)) was conducted by using the secondary electron imaging at an accelerating voltage of 10 kV. The chemical composition was analyzed by energy dispersive spectrometer (EDS) that attached to the SEM. The backscattered electron image (inset of Fig. 2(b)) was taken with an accelerating voltage of 15 kV. The size of the crystalline and surface roughness were obtained by an atomic force microscopy (AFM). The surface diffraction pattern was obtained by an in-situ high-pressure reflection high-energy electron diffraction (HP-RHEED) with a voltage of 30 kV after growth. Transport properties were measured by a physical property measurement system (PPMS, Quantum Design) with a magnetic field up to 9 T.
@@ III.1 Structural characterizations
Fig. 1(a) shows the wide-range XRD 2- scans for the metallic La3Ni2O7-δ thin films (sample SH2 and SH3). All the (00) ( is even) peaks of La3Ni2O7-δ appeared without any obvious impurity peaks, suggesting the -axis orientation of the films without any visible impurity phases. Fig. 1(b)-(e) show the enlarged XRD 2- scans of as-grown films around the La3Ni2O7-δ (006) peak (b and c) and (0010) peak (d and e). The dashed lines in grey, light red, and light green represent the peak positions of , , and La4Ni3O10 phases, respectively. The XRD 2- measurements indicate that all of the thin films display the dominant La3Ni2O7-δ phase. Samples in Fig. 1(b) and (d) were grown with lower laser energy (laser fluence 1.52 ), and samples in Fig. 1(c) and (e) were grown with higher laser energy (laser fluence 2.07 ). The remarks in Fig. 1(b) and (c) mark the oxygen partial pressure for the growth of each sample. Sample SL4 and SH3 were post-annealed under oxygen partial pressure = Pa at for 24 and 48 h, respectively. Other samples were post-annealed under oxygen partial pressure = Pa at for 1 h. As shown in Fig. 1(g), the films grown with higher laser energy have better crystallinity than those with lower laser energy.
Fig. 2(a) shows the scanning electron microscopy secondary electron image of sample SH3. The surface of the film is relatively flat. As shown in Fig. 2(b), the distribution of all elements in the sample is relatively uniform. Since the film is very thin (12 nm), the concentration of nickel atoms is very low but still detectable (0.44). The contributions of other atomic concentrations mainly come from the LaAlO3 substrate. In addition, the oxygen content determined by the EDS measurement is not an accurate value. Fig. 2(c) shows the atomic force microscopy topography of sample SH3. The diameter of the grains is about 50100 nm as shown in the inset of Fig. 2(c), and the average roughness of the surface is about 0.32 nm. Fig. 2(d) shows the RHEED pattern of sample SH3. The clear diffraction streaks at high oxygen partial pressure of 35 Pa indicate a flat surface of the film.
@@ III.2 Transport measurements
Fig. 3(a) and (b) show the curves of La3Ni2O7-δ thin films grown with lower and higher laser energy, respectively. It is clear that the resistivity of the samples decreases with the increasing oxygen partial pressure for both lower and higher deposition energy. However, as shown in Fig. 3(a), samples synthesized with lower laser energy are weak insulators for all deposition oxygen partial pressures and annealing conditions. This may be attributed to the poor crystallinity of these samples, as shown in Fig. 1(f). As shown in Fig. 3(b), the samples synthesized with higher laser energy are insulating when the growth oxygen partial pressure is less than 30 Pa. With the oxygen partial pressure raised to 35 Pa, the La3Ni2O7-δ thin films demonstrate a clear metal-to-insulator transition as temperature decreases. Through long-time annealing at high oxygen partial pressure, the metal-to-insulator transition temperature can be lowered from 51.3 K (sample SH2) to 24.3 K (sample SH3). This suggests that the weak insulating behavior of La3Ni2O7-δ is related to oxygen vacancies, consistent with previous studies [39, 40, 41]. It is noteworthy that all the curves in the whole measurement temperature region do not show the DW-like or SDW-like anomalies as seen in bulk samples. The possible transition may be suppressed by the substrate strain, as that in the bulks is suppressed by physical pressure [15, 17, 32]. The inset of Fig. 3(b) shows the resistivity behavior of sample SH2 and SH3 at low temperatures on semi-log plots. A clear logarithmic increase of the resistivity with decreasing temperature is observed. This may correlate with Kondo effect or weak localization in a Quasi-two-dimensional (2D) system, as will be addressed below.
Compared with the rocking curves in Fig. 1(f), the full-width-at-half-maximum on the La3Ni2O7-δ (006) peaks of samples grown at lower laser energy are generally wider than those grown at higher energy. And the rocking curve peaks of metallic samples are much sharper than that of the weak insulating samples, indicating the crystallinity might be important to gain metallic La3Ni2O7-δ thin films.
@@
To further investigate the electronic scattering and the information of the Fermi surface, we systematically measured the magnetoresistance of the metallic La3Ni2O7-δ thin film. The magnetoresistivity is defined as , where is the longitudinal resistivity in a magnetic field and is that at a zero-field. As shown in Fig. 5(a), the magnetoresistance above 50 K is positive but very weak (less than 0.1), indicating the absence of multiband effect. The negative magnetoresistance at temperatures below 40 K in Fig. 5(a) indicates the delocalization effect in a magnetic field, and it decreases with increasing temperature. The MR above 60 K is positive, and with the increasing temperature, it increases in the temperature region of 60140 K and decreases slightly at temperatures above 160 K. Fig. 5(b) shows the temperature dependence of the magnetoresistance at 9 T. It can be observed that the negative magnetoresistance at low temperatures exhibits a rapid decay with increasing temperature, while the positive magnetoresistance at high temperatures is notably weak and exhibits minimal variation with temperature. The inset of Fig. 5(b) shows the curve at 0 T and 9 T. The magnetic field exerts a slight suppression on the upturn of resistivity at low temperatures in a magnetic field of 9 T. According to Kohler’s law [46], if there is only one isotropic scattering time in the transport properties of a system, the magnetoresistance , where and represent some unknown functions (assuming that the scattering rate ). Fig. 5(c) shows the positive MR plotted by following Kohler’s law, One can see that Kohler’s law is obeyed at temperatures above about 180 K, disobeyed at temperatures below 160 K. As mentioned before, the variation of DOS with temperature due to correlation effect can make the basic requirement of Kohler’s law that invalid, which may cause the failure of Kohler’s law for describing the magnetoresistance.
@@ III.3 Analysis for the resistivity and negative magnetoresistance at low temperatures
Where is the residual resistivity, the term represents the contribution of Fermi liquid behavior, is the Kondo scattering term that contributes to the logarithmic rise behavior at low temperatures, and it can be expressed by Hamann formula [47]
Where is a temperature-independent constant, is the Kondo temperature, is the spin of the magnetic impurities. The fit by eq. 1 of the low-temperature resistivity is shown as the solid line in Fig. 6, with K and .
@@
Fig. 7 shows as a function of . The curves for K overlap each other. This means that the Kondo temperature is between 67.5 K, consistent with the fit result of the curve that K in Fig. 6. The lines in Fig. 7 represent the Brillouin function fits (eq. 3) for K. It’s clear that the curves for K can be well fitted by the Brillouin function with , which means that the local magnetic moment of the impurity is .
@@ IV SUMMARY
In summary, we have successfully synthesized (00)-oriented La3Ni2O7-δ thin films on LaAlO3 single crystal substrates by PLD technique. We have obtained samples with weak insulating and metallic behavior by adjusting the laser energy and oxygen partial pressure during growth. The rocking curves show that the crystallinity of the metallic samples is significantly better than that of the weak insulating samples. Moreover, the curves of all the La3Ni2O7-δ thin films in the temperature region of 2300 K do not show any anomalies corresponding to the spin or the charge density waves that have been observed in bulks. This may be attributed to the influence of the compressive strain provided by the substrate. Hall resistance shows a linear field dependence with the dominant hole charge carriers, but the Hall coefficient exhibits strong temperature dependence. The MR above about 50 K is positive but very weak, indicating the absence of multiband effect. The Hall angle shows that the scattering rate is close to , but the scattering rate derived from resistivity shows the temperature dependence with an exponent of about 1.5, which is explained by the involvement of correlation effect in the La3Ni2O7-δ thin films. Moreover, the magnetoresistance at low temperatures is negative, showing the delocalization effect. Detailed analysis of the magnetoresistance suggests that the delocalization effect can be attributed to the Kondo-like scattering, rather than the Anderson weak localization. The Kondo-like scattering suggests a Hund’s coupling between the orbital with the orbital.
@@ Acknowledgements.
We thank B. Yang and Y. Z. Tian for assistance with XRD and X-ray reflectivity analysis. We acknowledge financial support from the National Key Research and Development Program of China (No. 2022YFA1403201), National Natural Science Foundation of China (Nos. 11927809, 12061131001).References
@@
- [15] Y. Zhang, D. Su, Y. Huang, Z. Shan, H. Sun, M. Huo, K. Ye, J. Zhang, Z. Yang, Y. Xu, et al., Nat. Phys. (????).
@@
- Zhang et al. [2024] M. Zhang, C. Pei, X. Du, W. Hu, Y. Cao, Q. Wang, J. Wu, Y. Li, H. Liu, C. Wen, et al. (2024), eprint arXiv:2311.07423.

### v2

@@ Abstract
By using the pulsed-laser-ablation technique, we have successfully grown the La3Ni2O7-δ thin films with -axis orientation perpendicular to the film surface. X-ray diffraction shows that the (00l) peaks can be well indexed to the La3Ni2O7-δ phase. Resistive measurements show that the samples can be tuned from weak insulating to metallic behavior through adjusting the growth conditions. Surprisingly, all curves of in the temperature region of 2300 K do not show the anomalies corresponding to either the spin density wave or the charge density wave orders as seen in bulk samples. Hall effect measurements show a linear field dependence with the dominant hole charge carriers, but the Hall coefficient exhibits strong temperature dependence. The magnetoresistance above about 50 K is positive but very weak, indicating a weakened or absence of multiband effect. However, a negative magnetoresistance is observed at low temperatures, which shows the delocalization effect by magnetic field. Detailed analysis on the magnetoresistance suggests that the delocalization effect at low temperatures is due to the Kondo-like effect, rather than the Anderson weak localization. Our transport results suggest that, the electronic conduction is fulfilled by the orbital with holes as the dominant charge carriers, while the interaction through Hund’s coupling with the localized orbital plays an important role in the charge dynamics.
@@ II EXPERIMENTAL DETAILS
The La3Ni2O7-δ thin films were grown on single-crystal (00)-oriented LaAlO3 substrates by the technique of pulsed-laser deposition, with substrate temperature at and laser frequency of 2 Hz for all cases. Prior to the deposition, a polycrystalline La3Ni2O7-δ target was prepared by the sol-gel method [38]. During the growth, the laser spot area of 4.68 was fixed, and two laser fluence of 1.52 (labelled as SL) and 2.07 (labelled as SH) were applied. And the oxygen partial pressure during the growth was 2035 Pa. After growth, the samples were cooled down at a cooling rate of to under oxygen partial pressure =58 Pa and post-annealed in-situ at for about 148 h, followed by the same cooling procedure to room temperature. Then they were taken out for the subsequent measurements. XRD and X-ray reflectivity (XRR) measurements were performed by using a Bruker D8 Advanced diffractometer with Cu radiation. The reciprocal space map (RSM) is measured by a Bruker D8 discover diffractometer. The morphology characterization obtained by a scanning electron microscope (SEM) (Fig. 2(a)) was conducted by using the secondary electron imaging at an accelerating voltage of 10 kV. The size of the crystalline and surface roughness were obtained by an atomic force microscopy (AFM). The surface diffraction pattern was obtained by an in-situ high-pressure reflection high-energy electron diffraction (HP-RHEED) with a voltage of 30 kV after growth. Transport properties were measured by a physical property measurement system (PPMS, Quantum Design) with a magnetic field up to 9 T.
@@ III.1 Structural characterizations
Fig. 1(a) shows the wide-range XRD 2- scans for the metallic La3Ni2O7-δ thin films (sample SH2 and SH3). All the (00) ( is even) peaks of La3Ni2O7-δ appeared without any obvious impurity peaks, suggesting the -axis orientation of the films without any visible impurity phases. Fig. 1(b)-(e) show the enlarged XRD 2- scans of as-grown films around the La3Ni2O7-δ (006) peak (b and c) and (0010) peak (d and e). The dashed lines in grey, light red, and light green represent the peak positions of , , and La4Ni3O10 phases, respectively. The XRD 2- measurements indicate that all of the thin films display the dominant La3Ni2O7-δ phase. Samples in Fig. 1(b) and (d) were grown with lower laser energy (laser fluence 1.52 ), and samples in Fig. 1(c) and (e) were grown with higher laser energy (laser fluence 2.07 ). The remarks in Fig. 1(b) and (c) mark the oxygen partial pressure for the growth of each sample. Sample SL4 and SH3 were post-annealed under oxygen partial pressure = Pa at for 24 and 48 h, respectively. Other samples were post-annealed under oxygen partial pressure = Pa at for 1 h. As shown in Fig. 1(f), the films grown with higher laser energy have better crystallinity than those with lower laser energy. Fig. 1(g) shows the X-ray reflectivity of sample SH3 with corresponding fit shown in a red dashed line, which yields a film thickness of about 12.056 nm. The thicknesses of samples SH1 and SH2 are around 20 nm. Other samples have also thickness around 12nm. Fig. 1(h) shows the reciprocal space map of sample SH3 around the (103) diffraction peak of the LaAlO3 substrate. The dashed line indicates that the La3Ni2O7-δ thin film has a in-plane constrain by the substrate. The diffraction peak broadening on the left-hand side of the La3Ni2O7-δ thin film is slightly stronger than that on the right-hand side, which may be due to a slight relaxation in the surface layer. The -axis constant is calculated to be between 3.792 Å and 3.807 Å, smaller than that of the bulk samples. And the -axis constant is around 20.722 Å, slightly larger than that of the bulk sample due to the in-plain compressive strain.
Fig. 2(a) shows the scanning electron microscopy secondary electron image of sample SH3. The surface of the film is relatively flat. Fig. 2(b) shows the atomic force microscopy topography of sample SH3. The diameter of the grains is about 50100 nm as shown in the inset of Fig. 2(b), and the average roughness of the surface is about 0.32 nm. Fig. 2(c) shows the RHEED pattern of sample SH3. The clear diffraction streaks at high oxygen partial pressure of 35 Pa indicate a flat surface of the film.
@@ III.2 Transport measurements
Fig. 3(a) and (b) show the curves of La3Ni2O7-δ thin films grown with lower and higher laser energy, respectively. It is clear that the resistivity of the samples decreases with the increasing oxygen partial pressure for both lower and higher deposition energy. However, as shown in Fig. 3(a), samples synthesized with lower laser energy are weak insulators for all deposition oxygen partial pressures and annealing conditions. We also tried growth with higher growth partial pressures of oxygen (40 Pa) and with lower laser fluence. But the dominant phase of the thin film becomes La4Ni3O10 instead of the metallic La3Ni2O7-δ. This may be the reason for the poor crystallinity of the La3Ni2O7-δ in the these samples, as shown in Fig. 1(f). As shown in Fig. 3(b), the samples synthesized with higher laser energy are insulating when the growth oxygen partial pressure is less than 30 Pa. With the oxygen partial pressure raised to 35 Pa, the La3Ni2O7-δ thin films demonstrate a clear metal-to-insulator transition as temperature decreases. Through long-time annealing at high oxygen partial pressure, the metal-to-insulator transition temperature can be lowered from 51.3 K (sample SH2) to 24.3 K (sample SH3). This may suggests that the weak insulating behavior of La3Ni2O7-δ is related to oxygen vacancies, consistent with previous studies [39, 40, 41, 42]. It is noteworthy that all the curves in the whole measurement temperature region do not show the DW-like or SDW-like anomalies as seen in bulk samples. The possible transition may be suppressed by the substrate strain, as that in the bulks is suppressed by physical pressure [15, 17, 32]. The inset of Fig. 3(b) shows the resistivity behavior of sample SH2 and SH3 at low temperatures on semi-log plots. A clear logarithmic increase of the resistivity with decreasing temperature is observed. This may correlate with Kondo effect or weak localization in a Quasi-two-dimensional (2D) system, as will be addressed below.
We also performed some ex-situ oxygen annealing on these samples. We annealed these samples in an atmosphere of oxygen using a thermal gravimetric analyzer. However, the samples became more insulating. This may be due to the heat-treatments which worsens the film crystallinity. We then tried to process the insulating samples at a high oxygen pressure of 8 MPa, but even after annealing for a long time (60 h) at such a high oxygen pressure, they still could not become metallic, only a structural phase transition appeared at around 250 K. Unlike the bulk samples, where oxygen treatment can change the samples from insulating to metallic, for the film sample, it depends more on the growth process rather than on the oxygen treatment concerning whether it is metallic or insulating . Compared with the rocking curves in Fig. 1(f), the full-width-at-half-maximum on the La3Ni2O7-δ (006) peaks of samples grown at lower laser energy are generally wider than those grown at higher energy. And the rocking curve peaks of metallic samples are much sharper than that of the weak insulating samples, indicating the crystallinity might be important to gain metallic La3Ni2O7-δ thin films.
@@
To further investigate the electronic scattering and the information of the Fermi surface, we systematically measured the magnetoresistance of the metallic La3Ni2O7-δ thin film. The magnetoresistivity is defined as , where is the longitudinal resistivity in a magnetic field and is that at a zero-field. As shown in Fig. 5(a), the magnetoresistance above 50 K is positive but very weak (less than 0.1), indicating a weakened or absence of multiband effect. Thus, we think that the strong temperature dependence of the Hall coefficient should be partially contributed by the multiband effect which becomes weaker in high temperature region (due to temperature enhanced interband scattering), and partially due to the correlation effect, because the effective DOS at Fermi energy is temperature-dependent and becomes depleted with decreasing temperature. The negative magnetoresistance at temperatures below 40 K in Fig. 5(a) indicates the delocalization effect in a magnetic field, and it decreases with increasing temperature. The MR above 60 K is positive, and with the increasing temperature, it increases in the temperature region of 60140 K and decreases slightly at temperatures above 160 K. Fig. 5(b) shows the temperature dependence of the magnetoresistance at 9 T. It can be observed that the negative magnetoresistance at low temperatures exhibits a rapid decay with increasing temperature, while the positive magnetoresistance at high temperatures is notably weak and exhibits minimal variation with temperature. The inset of Fig. 5(b) shows the curve at 0 T and 9 T. The magnetic field exerts a slight suppression on the upturn of resistivity at low temperatures in a magnetic field of 9 T. According to Kohler’s law [47], if there is only one isotropic scattering time in the transport properties of a system, the magnetoresistance , where and represent some unknown functions (assuming that the scattering rate ). Fig. 5(c) shows the positive MR plotted by following Kohler’s law, One can see that Kohler’s law is obeyed at temperatures above about 180 K, disobeyed at temperatures below 160 K. As mentioned before, the variation of DOS with temperature due to correlation effect can make the basic requirement of Kohler’s law that invalid, which may cause the failure of Kohler’s law for describing the magnetoresistance.
@@ III.3 Analysis for the resistivity and negative magnetoresistance at low temperatures
Where is the residual resistivity, is the Kondo scattering term that contributes to the logarithmic rise behavior at low temperatures, and it can be expressed by Hamann formula [48]
Where is a temperature-independent constant, is the Kondo temperature, is the spin of the magnetic impurities. In order to improve the reliability of fitting results, we first fit the curve in the temperature region of 5080 K with (since the negative magnetoresistance disappears at above 50 K and the temperature region for the Kondo fit is chosen for temperatures between 2 to 80 K), then put the value of the resulting into the Kondo fitting formula and make sure that the values of and obtained from the Kondo fit are close to those obtained from the fit for 5080 K. The curve fitted by eq. 2 of the low-temperature resistivity is shown as the green solid line in Fig. 6, with K and .
As shown by the red solid line in Fig. 6, it gives K and , being close to the fitting results by eq. 2. Since it is a fit with multiple fitting parameters, the final fitting result may subject to some errors, but the Kondo-like scattering picture seems to be substantiated by the general behavior of the integrated data system.
@@
Fig. 7 shows as a function of . The curves for K overlap each other. This means that the Kondo temperature is between 67.5 K, slightly smaller than the value obtained from the fits of the curve. The lines in Fig. 7 represent the Brillouin function fits (eq. 4) for K. The raw data for K lies between the fitting curves of the Brillouin function with and , which means that is likely to be bewteen 0.5 and 1, slightly larger than that obtained from fits of the curve. The difference is probably mainly due to the large errors of the fits for the curve, since there are too many parameters. As metioned in the previous references [16, 42, 54, 55, 56, 57], the Ni electrons are localized forming the localized magnetic moment, the Ni electrons are itinerant. The interaction between electrons of these two orbitals may give rise to the Kondo effect. On the other hand, the fit of the magnetoresistance at low temperatures support that the spin of the magnetic impurities is close to 1. Since Ni is the only magnetic element in the La3Ni2O7-δ thin film and its divalent ion has a spin , there may be a small amount of impurities in the sample, causing the Kondo effect.
@@ IV SUMMARY
In summary, we have successfully synthesized (00)-oriented La3Ni2O7-δ thin films on LaAlO3 single crystal substrates by PLD technique. We have obtained samples with weak insulating and metallic behavior by adjusting the laser energy and oxygen partial pressure during growth. The rocking curves show that the crystallinity of the metallic samples is significantly better than that of the weak insulating samples. Moreover, the curves of all the La3Ni2O7-δ thin films in the temperature region of 2300 K do not show any anomalies corresponding to the spin or the charge density waves that have been observed in bulks. This may be attributed to the influence of the compressive strain provided by the substrate. Hall resistance shows a linear field dependence with the dominant hole charge carriers, but the Hall coefficient exhibits strong temperature dependence. The MR above about 50 K is positive but very weak, indicating a weakened or absence of multiband effect. The Hall angle shows that the scattering rate is close to , but the scattering rate derived from resistivity shows the temperature dependence with an exponent of about 1.5, which is explained by the involvement of correlation effect in the La3Ni2O7-δ thin films. Moreover, the magnetoresistance at low temperatures is negative, showing the delocalization effect. Detailed analysis of the magnetoresistance suggests that the delocalization effect can be attributed to the Kondo-like scattering, rather than the Anderson weak localization. The Kondo-like scattering suggests a Hund’s coupling between the orbital with the orbital.
@@ Acknowledgements.
We thank B. Yang and Y. Z. Tian for assistance with XRD and X-ray reflectivity analysis. Thank B. Hao for help with the RSM measurement and analysis. We acknowledge financial support from the National Key Research and Development Program of China (No. 2022YFA1403201), National Natural Science Foundation of China (Nos. 11927809, 12061131001).References
@@
- Zhang et al. [2024a] Y. Zhang, D. Su, Y. Huang, Z. Shan, H. Sun, M. Huo, K. Ye, J. Zhang, Z. Yang, Y. Xu, et al., Nat. Phys pp. 1–5 (2024a).
@@
- Zhang et al. [2024b] M. Zhang, C. Pei, X. Du, W. Hu, Y. Cao, Q. Wang, J. Wu, Y. Li, H. Liu, C. Wen, et al. (2024b), eprint arXiv:2311.07423.
@@
- Taniguchi et al. [1995] S. Taniguchi, T. Nishikawa, Y. Yasui, Y. Kobayashi, J. Takeda, S.-i. Shamoto, and M. Sato, J. Phys. Soc. Jpn. 64, 1644 (1995).
@@
- Goldhaber-Gordon et al. [1998] D. Goldhaber-Gordon, J. Göres, M. Kastner, H. Shtrikman, D. Mahalu, and U. Meirav, Phys. Rev. Lett 81, 5225 (1998).
- Lee et al. [2011] M. Lee, J. Williams, S. Zhang, C. D. Frisbie, and D. Goldhaber-Gordon, Phys. Rev. lett 107, 256601 (2011).
@@
- Qu et al. [2024] X.-Z. Qu, D.-W. Qu, J. Chen, C. Wu, F. Yang, W. Li, and G. Su, Phys. Rev. Lett 132, 036502 (2024).
- Ouyang et al. [2024] Z. Ouyang, J.-M. Wang, J.-X. Wang, R.-Q. He, L. Huang, and Z.-Y. Lu, Phys. Rev. B 109, 115114 (2024).
- Cao and Yang [2024] Y. Cao and Y.-f. Yang, Phys. Rev. B 109, L081105 (2024).
- Oh and Zhang [2023] H. Oh and Y.-H. Zhang, Phys. Rev. B 108, 174511 (2023).